OpenStudy (anonymous):

which of the following is not a Pythagorean triple? A. 3-4-5 B. 5-12-13 C. 7-24-25 D. 8-15-18

OpenStudy (jennychan12):

test each pair. a^2+b^2 = c^2

OpenStudy (anonymous):

D. c = 17

OpenStudy (anonymous):

u sure its D?

OpenStudy (anonymous):

\[\sqrt{8{}^{\wedge}2+15{}^{\wedge}2}=\sqrt{64+225}=\sqrt{289}=17 \]